Two moles of nitrogen gas at 25°C, confined within a cylinder by a piston maintaining a constant pressure of 1 atm, is heated with 5.30 kJ of energy. Assume all the energy is used to do work of expansion of the gas at 1 atm. What will be the final temperature of the gas? Recall ∆H = ∆E + P∆V and watch your units!
Answer #1

solution - *- Cose-0 -*-*- initicel Temperature I=85°C I or T = 25+273 | T = 898° Mlumber of moler - 1:37 initial pressure (PAf=o 7 0 Cose - 7 -&- Assume - Mitrogen gas is ideal gas- and final Temporature of gas is (I). → ideal-gas lover PV=nRT sinceI cuork (u) = par l criven AE=o, & pav=nRl Toloni out in the above equation - AH о+nRAT or AH = RR (Tartig But the valid of ATo = 29 8kt 55008 ampll x 8:319 $ mol. K T, = 998€ + 5300 K 2x 8.314 To = (298+318.79) K | Tg = 616.74k) To = 616074973 or |